# Kalman Filter Enhanced GRPO for Reinforcement Learning-Based Language Model Reasoning

### v2

@@ Abstract
Reward baseline is important for Reinforcement Learning (RL) algorithms to reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) is proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to inaccurate advantage estimates in environments with highly noisy rewards, potentially introducing bias. In this work, we propose a model, called Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), by using lightweight Kalman filtering to dynamically estimate the latent reward mean and variance. This filtering technique replaces the naive batch mean baseline, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through accuracy and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 1 Introduction
In reinforcement learning algorithms that directly optimize the policy via gradient estimates, including Policy Gradient (PG) (sutton1999policy, ), Actor-Critic (AC) (mnih2016asynchronous, ) and Proximal Policy Optimization (PPO) (schulman2017proximal, ), variance reduction of policy gradient estimate is critical for stable and sample-efficient training. A common approach to estimate the advantage of actions is to subtract a baseline, which can be interpreted as the expectation of returns by taking all actions for the current state, from the observed returns. By doing so, the model is ideally expected to approximate the expected value of the return under the current policy. In many popular algorithms, such as AC and PPO, this baseline is often computed via an additionally maintained value model. However, the value model introduces further computational costs, especially for large model training.
To overcome this issue, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) is proposed to simplify the prediction of reward baselines by taking the mean reward over a batch or group of trajectories. But this approach can introduce bias if inaccurate action advantage was estimated when the reward is highly non-stationary and sparse. This bias may lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we take one step ahead by treating the observed rewards as noisy measurements of a latent reward signal and use the Kalman filter to model the underlying reward and noise trajectories. The resulting filtered baseline adapts over time, smoothing out noise while tracking underlying values and uncertainties in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with the existing GRPO algorithm and requires no modification to the policy networks. It provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach and show how it improves training stability. In summary, this paper makes the following contributions:
@@
2 Related Work
2.1 Deep Reinforcement Learning
Reinforcement learning (RL) algorithms do not require strong supervision signals as supervised learning does, and they open another door for surpassing human performance. RL methods can be broadly divided into model-based and model-free methods. Model-based methods explicitly learn a model of the environment dynamics and use it for planning or policy improvement. These approaches can be sample-efficient but often suffer from model bias. In contrast, model-free methods directly learn a policy or value function from interaction with the environment without modeling the environment dynamics. Despite a higher sample complexity, model-free methods remain dominant in many practical applications due to their relative simplicity and robustness.
Among model-free approaches, value-based methods learn to approximate the expected return of actions under a given policy. The Deep Q-Network (DQN) (mnih2013playing, ) is a foundational algorithm that learns a Q-function using temporal difference learning following the Bellman Equation. Several extensions improve its stability and performance, including Double DQN (van2016deep, ) to address Q-value overestimation; Dueling DQN (wang2016dueling, ), which decomposes Q-values into state-value and advantage terms.
Policy-based methods directly optimize the policy using gradient estimates. The Policy Gradient (PG) method (sutton1999policy, ) uses sampled trajectories to compute the gradient of the expected return. Trust Region Policy Optimization (TRPO) (schulman2015trust, ) introduces a trust region constraint to ensure stable policy updates, while the Proximal Policy Optimization Algorithm (PPO) (schulman2017proximal, ) simplifies this with a clipped surrogate objective.
Combining both policy and value learning, actor-critic methods were proposed. This family includes Advantage Actor-Critic (A2C), which uses a critic to estimate the advantage function, and its asynchronous version Asynchronous Advantage Actor-Critic (A3C) (mnih2016asynchronous, ), which improves stability and exploration with multi-thread training. PPO can also be viewed as an actor-critic method with a clipped objective. In the continuous action setting, Deep Deterministic Policy Gradient (DDPG) (lillicrap2015continuous, ) learns a deterministic actor and a critic. Twin Delayed Deep Deterministic policy gradient (TD3) (fujimoto2018addressing, ) improves DDPG by addressing function overestimation and policy update delays. Soft Actor-Critic (SAC) (haarnoja2018soft, ) introduces entropy regularization to encourage exploration and improve stability.
These Reinforcement Learning algorithms can perform well for sequential action control, e.g. playing Atari games (van2016deep, ), MoJoCo (todorov2012mujoco, ), or other different simulations (wang2021multi, ; wang2020soft, ). However, they are not specifically designed for large language models (LLMs). Reducing the variance of policy gradient estimates is important for achieving stable and sample-efficient training. One widely used method is to subtract a baseline from the observed returns. This baseline typically represents the expected return over all possible actions given the current state. With this subtraction, the resulting action advantage estimation should align more closely with the expected return under the current policy. In commonly used algorithms like Actor-Critic and Proximal Policy Optimization, this baseline is usually obtained from an auxiliary value function model. Nevertheless, maintaining and training this additional model significantly increases computation, particularly when working with large-scale models.
2.2 Reinforcement Learning to Improve Reasoning for LLMs
Learning a reinforcement learning model from human feedback has been proposed by Christiano et al. (christiano2017deep, ) back in 2017. Recent advancements in aligning large language models (LLMs) with human preferences have predominantly utilized reinforcement learning from human feedback (RLHF) (ouyang2022training, ). RLHF typically involves training a reward model based on human-labeled preferences and subsequently fine-tuning the LLM using reinforcement learning algorithms such as Proximal Policy Optimization (PPO). This approach has been instrumental in enhancing models like ChatGPT (achiam2023gpt, ) and LLaMA (touvron2023llama, ).
However, RLHF presents challenges, including the complexity of reward model training and the instability of reinforcement learning procedures. To address these issues, alternative methods have been proposed. Direct Preference Optimization (DPO) (rafailov2023direct, ) eliminates the need for an explicit reward model by directly optimizing the policy to align with human preferences using a classification loss. This simplification leads to more stable and efficient training. Kahneman-Tversky Optimization (KTO) (ethayarajh2024kto, ) incorporates principles from prospect theory to better model human decision-making biases. By aligning the training objective with human utility functions, KTO aims to improve the alignment of LLM outputs with human preferences. Reinforced Self-Training (ReST) (gulcehre2023reinforced, ) adopts an offline reinforcement learning approach, generating training data from the model’s own outputs and refining the policy without the need for online interaction or extensive human feedback. Rank Responses with Human Feedback (RRHF) (yuan2023rrhf, ) simplifies the alignment process by ranking multiple model-generated responses and optimizing the model to prefer higher-ranked outputs, reducing reliance on complex reward models. Reinforcement Learning from AI Feedback (RLAIF) (lee2023rlaif, ) replaces human feedback with evaluations from a separate LLM, enabling scalable alignment without the need for human-labeled data. Self-Play Fine-Tuning (SPIN) (chen2024self, ) leverages self-play mechanisms, allowing the model to generate and learn from its own data, progressively improving performance without additional human supervision. Group Relative Policy Optimization (GRPO) introduces a group-based approach to policy optimization, where the model’s outputs are evaluated relative to a group baseline.
These methods represent a focus towards more efficient and scalable approaches for aligning LLMs with human preferences, reducing dependence on extensive human feedback and complex reinforcement learning procedures. But there is limited exploration from the estimation of a more accurate advantage that can lift up model performance.
Therefore, in this paper, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO), which replaces the group-mean-based advantage estimation in GRPO with a Kalman-filter-based one. By treating observed rewards as noisy measurements of a latent reward signal, the filter adaptively estimates a baseline that tracks underlying uncertainties. With non-parametric one-dimensional Kalman filter, we can have more accurate advantage estimation by filtering out the underlying reward baseline and uncertainty with minimal computational cost.
3 Methodology
@@ 3.1 Problem Setting
We consider reinforcement learning where an agent interacts with an environment modeled as a Markov Decision Process (MDP). At a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, and receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantage function to reduce the variance of gradient updates.
Group Relative Policy Optimization (GRPO) framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting the group sample mean. While this reduces variance, the group mean may be an inaccurate estimate of the expected reward for individual outputs, particularly when the reward distribution is noisy. This can lead to unstable or suboptimal updates during training.
@@ 3.2 Kalman-Filtered Advantage Estimation
We propose to replace the naive group mean baseline with a Kalman filter that maintains an online estimate of the reward distribution’s mean and variance. Each observed return $r_{i}$ is treated as a noisy measurement of the latent reward signal $x_{i}$ which we cannot obtain directly.
The Kalman filter maintains a posterior estimate of the reward baseline $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of $x_{i}$ given observations up to step $i-1$, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $\varepsilon$ is a small constant for numerical stability. This formulation has two benefits: it adaptively centers the reward using a filtered baseline, and it normalizes the advantage by the estimated uncertainty, which helps stabilize the scale of policy updates.
The filtered advantage estimate $A_{i}$ can be directly used in place of existing advantage estimates for models within the policy gradient family. In our implementation, we integrate it with GRPO by replacing the typical advantage computation with the Kalman-filtered version described above. Worth to be noted, this method does not require training a separate value network for baseline estimation, nor does it add new optimization objectives. The filter state is updated online during each policy update step, and incurs negligible additional computational cost.
@@ 3.3 KRPO Algorithm
The KRPO procedure is fully differentiable and can be implemented efficiently in the GRPO frameworks. After getting the group rewards, instead of getting the mean value from the group, we fetch the Kalman filtered baseline and the estimated uncertainty $P_{i|i}$.
By taking the posterior variance estimation $P_{i|i}$ into account, the advantage can be more effective to learn. That is to say, the “good” actions with large rewards can be better enhanced and the “poor” actions with small rewards can be further punished.
@@
4 Connections of Reward Mean and Kalman Filtering
The baseline used in GRPO model is the empirical mean of observed rewards. Specifically, for a group of returns $\{r_{i}\}_{i=1}^{n}$, the baseline is defined as
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i},$ (7)
This empirical mean can be interpreted as a special case of a Kalman filter with the following assumptions: (1) The process noise variance $Q\to 0$, indicating that the latent state (expected reward) is assumed to be constant; (2) The measurement noise variance $R$ is a constant; (3) no prior information for the initial state.
Specifically, consider the Kalman filter update equations as aforementioned Eq. 1 to Eq. 5. When $Q$ is constant and the system state does not change over time (i.e., the state transition matrix is an identity matrix and the process noise is zero), the Kalman filter update reduces to a weighted average of the observations, where the weights are determined by the prior variance and the observation noise variance. In practice, however, when the full batch of $\{r_{i}\}$ is available, computing the sample mean corresponds to performing a batch update using all data equally. This is equivalent to a Bayesian update with all measurements weighted equally, which can be seen as a limiting form of the Kalman filter without temporal dynamics.
This observation motivates the use of a Kalman filter, which allows slight noises when new rewards observed ($Q>0$), and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive and sequential-aware estimate of the reward baseline.
@@ 5 Experiments
5.1 Data
We use two datasets in our experiments, i.e., the Arithmetic Dataset and OpenMathInstruct Dataset, each of which is divided into three sub-datasets: easy, normal, and hard.
Arithmetic Dataset (tinygrpo2024, ). It contains 100,000 arithmetic problems involving addition, subtraction, multiplication, and division. Each problem is represented as a natural language prompt and the target is the correct answer as a string. The arithmetic expressions include up to 6-digit numbers and up to 6 terms. We first split the 100,000 problems into a training set and a testing set using the standard 80%/20% split. During both training and testing, we filter out problems whose natural language prompt exceeds 128 characters. Then, we categorize the problems into three sub-datasets based on the number of digits and terms: the easy set includes problems with at most 3-digit numbers and at most 3 terms; the normal set includes problems with at most 5-digit numbers and at most 5 terms; and the hard set includes problems with at most 6-digit numbers and at most 6 terms.
OpenMath-Instruct Dataset (toshniwal2024openmath, ). This dataset contains 7,500 math problems from various topics, including Algebra, Geometry, Prealgebra, Precalculus, Number Theory, Intermediate Algebra, and Counting & Probability. Each problem is labeled with one of five difficulty levels provided by the dataset. We split this dataset into a training set and a testing set using the same 80%/20% ratio. As with the first dataset, we filter out problems whose question text is longer than 128 characters during training and testing. We then group the problems into three sub-datasets based on difficulty level: the easy set includes problems with level $\leq$ 2; the normal set includes problems with level $\leq$ 3; and the hard set includes problems with level $\leq$ 5.
@@ 5.2 Implementation Details
We adopt a group size of 12 for both GRPO and our proposed KRPO models. The training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation managable, we adopt Llama-3.2-1B-Instruct as our base model, to individually train with GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and variance of the filter measurement noise $R$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, the training is conducted for 5,500 steps. For the reward design, following DeepSeek-R1-Zero model (guo2025deepseek, ) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48 Gigabyte of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.

### v3

@@ Abstract
The advantage function is a central concept in RL that helps reduce variance in policy gradient estimates. Recently, for language modeling, Group Relative Policy Optimization (GRPO) was proposed to compute the advantage for each output by subtracting the mean reward, as the baseline, for all outputs in the group. However, it can lead to high variance when the reward advantage is inaccurately predicted. In this work, we propose Kalman Filter Enhanced Group Relative Policy Optimization (KRPO) model, by using lightweight Kalman filtering to dynamically estimate the latent reward baseline and uncertainty. This filtering technique replaces the naive group mean, enabling more adaptive advantage normalization. Our method does not require additional learned parameters over GRPO. This approach offers a simple yet effective way to incorporate multiple outputs of GRPO into advantage estimation, improving policy optimization in settings where highly dynamic reward signals are difficult to model for language models. Through the accuracies and rewards obtained from math question answering and reasoning, we show that using a more adaptive advantage estimation model, KRPO can improve the stability and performance of GRPO. The code is available at https://github.com/billhhh/KRPO_LLMs_RL.
@@ 1 Introduction
In reinforcement learning algorithms that directly optimize the policy via gradient estimates, including Policy Gradient (PG) (Sutton et al., 1999), Actor-Critic (AC) (Mnih et al., 2016) and Proximal Policy Optimization (PPO) (Schulman et al., 2017), variance reduction of policy gradient estimate is critical for stable and sample-efficient training. A common approach to estimate the advantage of actions is to subtract a baseline (and divide by a variance), which can be interpreted as the expectation of returns by taking all sampled actions for the current state, from the observed returns. By doing so, the model is ideally expected to approximate the expected value of the return under the current policy. In many prevalent algorithms, such as AC or PPO, the baseline is often computed via an additional maintained value model. However, the value model introduces further computational costs, especially for large model training.
To overcome this issue, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is proposed to simplify the prediction of reward baselines by taking the mean reward over a group of trajectories. However, this approach can lead to high variance if an inaccurate advantage is estimated when the reward is highly nonstationary. This may lead to unstable updates and slower convergence, especially in settings where the reward signal evolves rapidly with high variance over time.
To address this issue, we propose a dynamic advantage estimation method based on a one-dimensional Kalman filter, named as Kalman Filter Enhanced Group Relative Policy Optimization (KRPO). Instead of relying on the group mean, we treat the observed rewards as noisy measurements of a latent reward baseline signal and use the Kalman filter to model the underlying reward and noise trajectories. The resulting filtered baselines and uncertainties adapt over time, smoothing out noise while tracking underlying reward baseline values in the environment. This enables more accurate and robust advantage estimation with minimal computational overhead. Our method integrates naturally with the existing GRPO algorithm and does not require modification of policy networks. Provides a principled way to incorporate uncertainty into advantage estimation without introducing additional learning components or tuning objectives. We evaluate our approach by comparing it to the standard group-mean approach and show how it improves training stability. In summary, this paper makes the following contributions:
@@
2 Kalman Filter Enhanced Group Relative Policy Optimization
@@ 2.1 Problem Setting
We consider reinforcement learning in which an agent interacts with an environment modeled as a Markov Decision Process (MDP). In a timestep, the agent observes a state $s$, selects an action $a$ according to a policy $\pi(a\mid s)$, and receives a reward $r$. The goal is to maximize the expected cumulative return. Policy gradient methods rely on estimating the advantages to reduce the variance of gradient updates. The Group Relative Policy Optimization (GRPO) framework is applied to large language models by sampling multiple responses (i.e., action trajectories) from the policy given the same input prompt. A shared baseline is computed as the average return across all sampled responses for the same prompt, and the advantage for each output is calculated by subtracting the group sample mean. Although this reduces variance, the group mean may result in inaccurate estimates of action advantages, particularly when the rewards are very noisy. This can lead to unstable or suboptimal updates during training.
@@ 2.2 Kalman-Filtered Advantage Estimation
We propose replacing the naive group mean baseline with a Kalman filter that maintains an online estimate of the expected reward and its variance to obtain a more accurate estimate of the advantages. Each observed return $r_{i}$ is treated as a noisy measurement of the latent expected reward signal $\hat{x}_{i}$, which we cannot obtain directly.
The Kalman filter maintains a posterior estimate of the reward baseline $\hat{x}_{i|i}$ and its associated uncertainty $P_{i|i}$, both updated recursively. Here, the subscripts follow the standard Kalman notation: $\hat{x}_{i|i-1}$ denotes the prior estimate of “genuine” reward baseline $\hat{x_{i}}$ given observations up to step i1i-1, while $\hat{x}_{i|i}$ is the posterior estimate after incorporating the observation $r_{i}$. Similarly, $P_{i|i-1}$ and $P_{i|i}$ denote the corresponding uncertainty (variance) estimates before and after observing $r_{i}$.
@@
where $\varepsilon$ is a small constant for numerical computation stability. This formulation has two benefits: it adaptively centers the reward using a filtered baseline, and it normalizes the advantage by the estimated uncertainty, which helps stabilize the scale of policy updates and get a more reasonable advantage estimate.
The obtained advantage estimate $A_{i}$ can be directly used in place of existing advantage estimates for models within the policy gradient family. In our implementation, we integrate it with GRPO by replacing the typical advantage computation with the Kalman-filtered version described above. This method does not require training a separate value network for advantage estimation, nor does it add new optimization objectives. The filter state is updated online during each policy update step, and incurs negligible additional computational cost.
@@ 2.3 KRPO Algorithm
The KRPO procedure is fully differentiable and can be implemented efficiently upon the GRPO frameworks. After getting the group rewards, instead of receiving the mean value from the group, we fetch the Kalman filtered baseline and estimated uncertainty $P_{i|i}$. By taking into account the posterior variance estimation $P_{i|i}$, the advantage can be more adaptively modeled.
@@
Worth noting, in the model training, the observations are completely randomly sampled in each iteration. Moreover, the order of samples in our implementation is arbitrary (e.g., they can be shuffled before processing), so no unintended temporal dependence is imposed on the data. It simply provides a recursive estimator with an accompanying measure of uncertainty. The ordering will thus not effect policy optimization.
Besides, each Kalman filter operation is applied to groups of rollouts. In the KRPO setting, although the reward values for individual samples are discrete, the grouped reward observations, which aggregate multiple rollouts, are not sparse and are much more continuous in distribution. According to the Central Limit Theorem (CLT, 2008), the sum of i.i.d. sampled rewards (with finite mean and finite variance) tends to follow a Gaussian distribution. This is consistent with the assumptions of our KRPO setting.
3 Connections of Group-mean and Kalman Filtering
The baseline computed in the GRPO model is the empirical mean of observed rewards. Specifically, for a group of returns $\{r_{i}\}_{i=1}^{n}$, the baseline is defined as
$\bar{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$ (7)
This empirical mean can be interpreted as a special case of a Kalman filter with the following assumptions: (1) The process noise variance $Q\to 0$, indicating that the latent state (expected reward) is assumed to be constant; (2) The measurement noise variance $R$ is a constant; (3) initial state is unknown, modeled with large uncertainty.
Specifically, consider the Kalman filter update equations as aforementioned Eq. 1 to Eq. 5. When $Q$ is constant and the system state does not change over time (i.e., the state transition matrix is an identity matrix and the process noise is zero), the Kalman filter update reduces to a weighted average of the observations, where the weights are determined by the prior variance and the observation noise variance. However, in practice, when the full batch of $\{r_{i}\}$ is available, computing the sample mean corresponds to performing a batch update using all data equally. This is equivalent to a Bayesian update with all measurements weighted equally, which can be seen as a limiting form of the Kalman filter without temporal dynamics.
This observation motivates the use of a Kalman filter, which accounts for both observation noise ($R>0$) and possible slight changes in the latent reward ($Q>0$), and dynamically adjusts the influence of new observations via the gain $K_{i}$. This leads to a more adaptive estimate of the reward advantage, making the Kalman filter a more general formulation compared to a simple average.
@@ 4.1 Implementation Details
We adopt four datasets in our experiments, including Arithmetic (Open-Thought, 2024), OpenMathInstruct (Toshniwal et al., 2024), MATH500 (Lightman et al., 2023), and AIME (Veeraboina, 2024). We keep the standard training/testing 80%/20% split ratio for each of the four datasets without model selection. Arithmetic and OpenMathInstruct are divided into three sub-datasets individually: easy, normal, and hard. More details about the datasets used in the paper can be found in the appendix. In the paper, we focus on math tasks. These tasks are suitable for experiments as they require the model to have adequate reasoning ability for resolving difficult math questions (even not easy for humans), and language model reasoning is the main focus of our KRPO paper. Moreover, math tasks have well-defined answers for reward design or introduce extra reward models, which is not the main emphasis of our paper.
We adopt a group size of 12 for both the GRPO and our proposed KRPO models. In our implementation, the Kalman filter state is reset for each prompt to avoid cross-prompt information leakage and ensure statistical independence across prompts. The training batch size is set to 16. The model is optimized using the Adam optimizer with an initial learning rate of $5\times 10^{-6}$. To make the computation manageable, we adopt Llama-3.2-1B-Instruct, Qwen2.5-0.5B-Instruct and Qwen2.5-1.5B-Instruct as our base models, to individually train with the GRPO algorithm and the proposed KRPO algorithm. By default, the Kullback–Leibler (KL) divergence term is weighted by a factor $\alpha$ of 0.01 during training; variance of the filter process noise $Q$ is set to 1e-5 and the variance of the filter measurement noise $R$ is set to 1e-2. For the Arithmetic dataset, the model is trained for 16,000 steps. For the OpenMathInstruct dataset, training is conducted for 5,500 steps. For both MATH500 and AIME datasets, we train the model for 10,000 steps. For the reward design, following the DeepSeek-R1-Zero model (Guo et al., 2025) to keep simplicity, we do not use a reward model and instead assign rewards as follows: if an answer is exactly equal to the ground-truth answer, the reward is 1.0; if a ground-truth answer is contained within a predicted answer, a 0.5 reward is assigned; otherwise, no reward will be returned. We also encourage the model to think reasonably through “<think></think>” tags. The random seed of experiments is set to 42. All experiments are conducted on one A6000 Nvidia GPU with 48GB of graphics memory. The proposed KRPO model and its compared models are trained and tested under the same setting to keep fair comparisons.